OpenStudy (anonymous):

Which inequality has the solution of a < - 4? a + 4 > 0 a – 1 < 3 a + 3 < -1 a – 4 > 0

OpenStudy (xapproachesinfinity):

okay so you a<-4 this can be written as a<-3-1 add 3 to both side a+3<-1 !

OpenStudy (xapproachesinfinity):

the other options you are given are different and don't satisfy this inequality

OpenStudy (xapproachesinfinity):

okay try and solve them and give me the answer

OpenStudy (anonymous):

Can you show me how?

OpenStudy (xapproachesinfinity):

no wrong! you cannot change the sign this subtraction not multiplication

OpenStudy (anonymous):

im not sure which one of you is right

OpenStudy (xapproachesinfinity):

a+4>0 sub 4 a>-4

OpenStudy (xapproachesinfinity):

My answer is right sir! xD

OpenStudy (anonymous):

so the correct answer is?

OpenStudy (xapproachesinfinity):

No my answer is completely right if you disagree show your proof

OpenStudy (xapproachesinfinity):

haha! well if wrong prove me wrong! that's very mathematic^_^

OpenStudy (xapproachesinfinity):

Okay no worries!

OpenStudy (anonymous):

can you help me with one more please

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The correct answer is the third one

OpenStudy (anonymous):

sure I'll try

OpenStudy (anonymous):

Evaluate 27 × a-2 for a = 3. one-third one-ninth 9 3

OpenStudy (anonymous):

thank you

OpenStudy (xapproachesinfinity):

Thanks meh!

OpenStudy (xapproachesinfinity):

you are welcome marissa^_^

OpenStudy (anonymous):

No problem now lets help our friend

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

is the x a multiplication sign or a variable

OpenStudy (anonymous):

multiplication sign

OpenStudy (anonymous):

I'm stumped let me ask for some other people to help you @ganeshie8 @xapproachesinfinity @Hero @JoannaBlackwelder @gswag98 @math92130

OpenStudy (xapproachesinfinity):

\(\large 27\times a^-2\)

OpenStudy (xapproachesinfinity):

that's a^-2

OpenStudy (xapproachesinfinity):

Now just put 3 in place of x

OpenStudy (joannablackwelder):

*In place of a.

OpenStudy (xapproachesinfinity):

27x(3)^-2=27/(3)^2

OpenStudy (xapproachesinfinity):

yeah in place of a i meant typo hehe thanks for the correction

OpenStudy (xapproachesinfinity):

so 27/9=3

OpenStudy (joannablackwelder):

No worries. You're doing great :)

OpenStudy (xapproachesinfinity):

thanks! ^_^

OpenStudy (anonymous):

so the answer would be?

OpenStudy (joannablackwelder):

27/9=3

OpenStudy (anonymous):

thank you so much

OpenStudy (xapproachesinfinity):

I gave you the answer hehe can't you see it lol

OpenStudy (xapproachesinfinity):

welcome!